OpenStudy (anonymous):

Use synthetic division to determine whether the number k is an upper or lower bound (as specified for the real zeros of the function f). k = 2; f(x) = 2x^3 + 3x^2 - 4x + 4; Lower bound?

OpenStudy (amistre64):

synthetic is a process, what about the process are you having an issue with?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

show me your work, or possibly ask a question about the process like how to start it, what to use, where did i go wrong ... etc

OpenStudy (anonymous):

I actually just got it. My coefficients got mixed up. They are all positive so it's in the upper bound. I'm sorry. But do you think you could help me solve a different question. I have no idea where to start with it.

OpenStudy (amistre64):

i might .. let see the question :)

OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = the quantity x minus seven divided by the quantity x plus three. and g(x) = quantity negative three x minus seven divided by quantity x minus one.

OpenStudy (amistre64):

do you recall what the notaiton f(4) means?

OpenStudy (amistre64):

also, if you can post a more mathical construction that would be helpful. words and math just dont mix

OpenStudy (anonymous):

The questions looks like f(x) = (x-7)/(x-3) and g(x) = (-3x - 7)/(x - 1)

OpenStudy (amistre64):

f(g(x)) is just when the x parts of f, are equal to g(x) f(g(x)) = (g(x)-7)/(g(x)-3) (g(x)-7-3+3)/(g(x)-3) 1 - 4/(g(x)-3) we have g(x) defined so plug it in and simplify to see if it reduces to x

OpenStudy (amistre64):

or work it into g(x) 1 - 4/[g(x)-3] = x 1 - x =4/[g(x)-3] g(x) - 3 =4/(1-x) g(x) = 4/(1-x) + 3 is this the definition of g(x) ?

OpenStudy (anonymous):

Yes?

OpenStudy (amistre64):

if you are unsure, then just work the rest of the algebra ...

OpenStudy (anonymous):

I have no idea what I'm doing :(

OpenStudy (amistre64):

which is why i asked to start with if you knew what the notation f(4) means, if you kow what this notation means then the rest is just process.

OpenStudy (anonymous):

I don't know what that is

OpenStudy (amistre64):

then you are not prepared to work this problem. spose we have a function f, defined in terms of x: then f(x) is the name of the function if we want to know what the value of f is when x=4, then we can express that by the notation f(x=4), or shorthanded to f(4) now, tell me what this means to you

OpenStudy (anonymous):

Nothing. It means absolutely nothing. I'm wasting your time. I'm just going to skip the questions. Thanks anyway.

OpenStudy (amistre64):

... wish i could have helped out better then, but good luck all the same :)